### v1

# VideoUFO: A Million-Scale User-Focused Dataset for Text-to-Video Generation

Abstract
Text-to-video generative models convert textual prompts into dynamic visual content, offering wide-ranging applications in film production, gaming, and education. However, their real-world performance often falls short of user expectations. One key reason is that these models have not been trained on videos related to some topics users want to create. In this paper, we propose VideoUFO, the first Video dataset specifically curated to align with Users’ FOcus in real-world scenarios. Beyond this, our VideoUFO also features: (1) minimal ($0.29\%$) overlap with existing video datasets, and (2) videos searched exclusively via YouTube’s official API under the Creative Commons license. These two attributes provide future researchers with greater freedom to broaden their training sources. The VideoUFO comprises over $1.09$ million video clips, each paired with both a brief and a detailed caption (description). Specifically, through clustering, we first identify $1,291$ user-focused topics from the million-scale real text-to-video prompt dataset, VidProM. Then, we use these topics to retrieve videos from YouTube, split the retrieved videos into clips, and generate both brief and detailed captions for each clip. After verifying the clips with specified topics, we are left with about $1.09$ million video clips. Our experiments reveal that (1) current $16$ text-to-video models do not achieve consistent performance across all user-focused topics; and (2) a simple model trained on VideoUFO outperforms others on worst-performing topics. The dataset is publicly available here under the CC BY 4.0 License.
footnotetext: *Corresponding author.
Figure 1: The glowing firefly: (a) generated by Sora [2] and (b) captured in a real video. The generated firefly is noticeably different from its real-life counterpart and thus unsatisfying. We attribute this primarily to a lack of exposure to such topics.
1 Introduction
Text-to-video generation aims to convert textual descriptions into dynamic visual content. Its applications are extensive and transformative, covering areas from creative media [4] and entertainment [43] to practical domains such as education [6], advertising [46], and assistance [14].
Despite their popularity and usefulness, current text-to-video models often fail to meet users’ expectations in real-world applications. For example, when we ask Sora [2] to generate a video using the prompt “A firefly is glowing on a grass’s leaf on a serene summer night”, it fails to capture the concept of a glowing firefly while successfully generating grass and a summer night, as shown in Fig. 1 (a). From the data perspective, we infer this is mainly because Sora [2] has not been trained on firefly-related topics, while it has been trained on grass and night. Furthermore, if Sora [2] has seen the video shown in Fig. 1 (b), it will understand what a glowing firefly should look like.
To this end, this paper presents VideoUFO, the first Video dataset curated specifically based on real-world Users’ FOcus in text-to-video generation. Such a dataset can help improve the alignment between text-to-video models and actual user needs. Specifically, we focus on: (1) curating VideoUFO by analyzing real users’ interests and scraping relevant videos; (2) comparing the VideoUFO with recent video datasets to highlight differences; and (3) demonstrating how our VideoUFO benefits video generation.
Figure 2: VideoUFO is the first dataset curated in alignment with real-world users’ focused topics for text-to-video generation. Specifically, the dataset comprises over $1.09$ million video clips spanning $1,291$ topics. Here, we select the top $20$ most popular topics for illustration. Researchers can use our VideoUFO to train or fine-tune their text-to-video generative models to better meet users’ needs.
The first dataset that aligns with real-world users’ focus in text-to-video generation. Our key idea is to analyze user-focused topics from user-provided prompts and then search for videos related to these topics. As shown in Fig. 2, the resulted VideoUFO comprises more than $\mathbf{1.09}$ million video clips spanning $\mathbf{1,291}$ user-focused topics. Specifically, we initiate by analyzing user-focused topics, which involves embedding all $1.67$ million user-provided prompts from VidProM [35], clustering these embeddings using K-means, and generating a topic for each cluster with GPT-4o [21], followed by combining similar topics. After obtaining these topics, we (1) search for these topics on YouTube, (2) segment the fetched videos into multiple semantically consistent short clips, (3) generate both brief and detailed captions for each clip, (4) filter out clips that do not contain the specific topics, and (5) assign each clip video quality scores that align with human perception. Note that although our current VideoUFO comprises about one million videos, it can be easily scaled up to ten million or more by sourcing additional videos for each topic. The VideoUFO can also be easily extended to the image-to-video domain by using the text and image prompts in TIP-I2V [34].
Differences between VideoUFO and other recent video datasets. Recently, several video datasets have been released, including OpenVid-1M [20], HD-VILA-100M [41], InternVid [38], Koala-36M [33], LVD-2M [40], MiraData [9], Panda-70M [5], VidGen-1M [29], and WebVid-10M [3]. While inheriting their excellent attributes – such as large scale, accurate captions, and high resolution – our VideoUFO explores some novel directions: (1) Guided by real user focus. Specifically, whereas recent datasets are typically gathered from open-domain sources, we concentrate on topics that are focused by text-to-video users. This feature enables text-to-video models trained on our dataset to better cater to users, while avoiding unnecessary expansion of the dataset and wasting resources. (2) Introducing new data. Although these recent papers claim to contribute new datasets, most of them primarily introduce a new data pipeline – that is, they reprocess HD-VILA-100M [41] and are subsets of it. While their contributions are useful and meaningful, in theory, a generative model that has already been fully fitted on HD-VILA-100M [41] would not gain new video knowledge from them. In contrast, we collect new data from YouTube, with only $0.29\%$ of the videos overlapping with existing datasets. (3) Data compliance. When curating VideoUFO, we retrieve videos using YouTube’s official API and select only those with a Creative Commons license. In contrast, most recent datasets do not explicitly address the regulatory compliance of their data collection process. This feature grants researchers greater flexibility in using our data.
Benchmarking current text-to-video models on user-focused topics and demonstrating the effectiveness of our VideoUFO. We present a new benchmark to quantify the performance of text-to-video models on user-focused topics. Specifically, our process involves: (1) selecting $10$ user-provided prompts per topic; (2) generating a video for each prompt; (3) using a multimodal large language model to describe each video; and (4) calculating the similarity between the generated descriptions and the original prompts. For each topic, we calculate the average similarity between the $10$ prompts and their corresponding descriptions. By sorting these averages, we can identify the worst-performing and best-performing topics for each model. Using this benchmark, we evaluate the performance of current text-to-video models and a newly trained text-to-video model on our VideoUFO. Experimental results indicate that (1) current $16$ text-to-video models have some poor-performing topics, and (2) our model achieves the highest similarities on worst-performing topics while maintaining performance on the best-performing ones.
In conclusion, our key contributions are as follows:
  1. 1.
    We present VideoUFO, the first video dataset curated based on the focus of real text-to-video users. This dataset comprises over $1.09$ million video clips spanning $1,291$ user-focused topics.
  2. 2.
    We compare VideoUFO with recent video datasets, highlighting their differences in both fundamental attributes and topics coverage, thereby emphasizing the necessity of our dataset. We also follow best practices in their curation processes to ensure the quality of our dataset.
  3. 3.
    We evaluate current text-to-video models on user-focused topics and observe that a simple model trained on our VideoUFO outperforms competing models on worst-performing topics.
2 Related Works
Text-to-Video Generation. The introduction of Sora [2] has sparked significant research interest in text-to-video generation. Commercial models such as Movie Gen [23], Veo [27], Kling [1], PixelDance [13], and Seaweed [13] demonstrate strong performance and are increasingly being integrated into various industries. Meanwhile, open-source models like HunyuanVideo [12], LTX-Video [15], CogVideoX [42], Mochi-1 [30], and Pyramidal [8] empower researchers to experiment, customize, and enhance existing frameworks. Although they are powerful, their training data determines the upper limit of the video quality they can generate. In this paper, we propose a dataset, VideoUFO, which has the potential to help these models better cater to users’ preferences in real-world applications.
Text-Video Datasets. A text-video dataset consists of video clips paired with corresponding textual descriptions or captions. Many text-video datasets, such as OpenVid-1M [20], HD-VILA-100M [41], InternVid [38], Koala-36M [33], LVD-2M [40], MiraData [9], Panda-70M [5], VidGen-1M [29], and WebVid-10M [3], have been proposed to advance multimodal understanding and generation. However, these datasets are not designed to align with the preferences of text-to-video users, creating a gap between training data topics and real-world applications. In contrast, our dataset VideoUFO is collected from a user-focus perspective. To better cater users, future researchers can fine-tune text-to-video models with the VideoUFO or integrate it with existing datasets to develop new models.
Preference Alignment in Diffusion Models. While diffusion models have demonstrated impressive capabilities, there is an increasing demand to ensure their outputs align with human preferences. Direct Preference Optimization (DPO) [25], originally proposed to align large language models (LLMs) with human preferences, can also be effectively applied to diffusion models. For example, DiffusionDPO [31] fine-tunes text-to-image models using human comparison data to improve image generation quality. Additionally, VideoDPO [17] introduces a comprehensive preference scoring system that evaluates both visual quality and semantic alignment to enhance text-to-video generation. Unlike other approaches that align text-to-image or text-to-video models with attributes such as aesthetics, motion, consistency, and visual appeal, our goal is to improve the performance of text-to-video models on real-world users’ focused topics.
[…]
Figure 4: Each data point in our VideoUFO includes a video clip, an ID, a topic, start and end times, a brief caption, and a detailed caption. Beyond that, we evaluate each clip with six different video quality scores from VBench [7].
3 Curating VideoUFO
Fig. 4 illustrates a data point from our million-scale VideoUFO, which comprises: (1) a video clip along with its ID; (2) the topic of the clip; (3) the start and end times within its original video; (4) the corresponding brief and detailed captions; and (5) the quality of the clip. The following provides an explanation of how we curate the VideoUFO.
Analyzing text-to-video users’ focused topics. We aim to analyze the topics that real users focus on or prefer when generating videos from texts. Here, we use VidProM [35] as it is the only publicly available, million-scale text-to-video prompt dataset written by real users. First, we embed all $1.67$ million prompts from VidProM into $384$-dimensional vectors using $\mathtt{SentenceTransformers}$ [26]. Next, we cluster these vectors with K-means. Note that here we pre-set the number of clusters to a relatively large value, i.e., $2,000$, and merge similar clusters in the next step. Finally, for each cluster, we ask GPT-4o [21] to conclude a topic ($1\sim 2$ words). The prompt is shown in the Supplementary (Section 8). After merging singular and plural forms, restoring verbs to their base forms, removing duplicates, and manually verifying the topics – removing overly broad ones such as ‘animation’, ‘scene’, ‘movement’, and ‘film’ – we finally obtained $\mathbf{1,291}$ topics. The semantic distribution of these topics is shown in Fig. 3, and the full list is available in the Supplementary (Section 12).
Figure 5: The statistical information of captions and video clips in the proposed VideoUFO dataset. The average word count for brief and detailed captions is $\mathbf{13.8}$ and $\mathbf{155.5}$, respectively, while the average clip duration is $\mathbf{12.6}$ seconds.
Collecting videos from YouTube. For each topic obtained, we use the official YouTube API to search for videos. Specifically, for each topic, we search for approximately $500$ videos by relevance, with each video meeting the following criteria: (1) it is shorter than $4$ minutes, (2) it has a resolution of $720$p or higher, and (3) it is licensed under Creative Commons. These requirements ensure that the collected videos are suitable and freely usable for training video generation models. In the end, we obtain $586,490$ videos from YouTube. We compare the YouTube IDs of these videos with those in existing datasets, including OpenVid-1M [20], HD-VILA-100M [41], InternVid [38], Koala-36M [33], LVD-2M [40], MiraData [9], Panda-70M [5], VidGen-1M [29], and WebVid-10M [3], and find that only $1,675$ IDs ($0.29\%$) are already present in these datasets. This suggests that our VideoUFO introduces novel information or knowledge, which can serve to expand the range of existing training sources.
Splitting videos and generating captions. After obtaining the videos, we segment them into multiple semantics-consistent short clips following the steps in curating Panda-70M [5]. The process includes shot boundary detection, stitching, and video splitting, producing a total of $\mathbf{3,181,873}$ clips. To facilitate the training of text-to-video models on our VideoUFO, we generate both brief and detailed captions for each video clip. For the brief captions, we utilize the video captioning model provided by Panda-70M [5]. For the detailed captions, we adopt the pipeline used in $\mathtt{Open}$-$\mathtt{Sora}$-$\mathtt{Plan}$ v1.3.0 [16], which employs $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ [32] for video annotation. In Fig. 5 (a) and (b), we present the statistical distribution of the lengths of brief and detailed captions, respectively.
Verifying clips. We notice that, although the videos are searched by topics, not every clip contains our intended topics. A straightforward solution is to use GPT-4o [21] to verify whether a given clip corresponds to a specific topic. However, while effective, this approach would be prohibitively expensive for verifying more than $3$ million video clips. As an alternative, we use the detailed caption (instead of the original clip) and feed it into GPT-4o mini (instead of GPT-4o) to verify whether it belongs to a specific topic. Since (1) GPT-4o mini and GPT-4o have similar capabilities in basic language understanding, and (2) the video understanding model provides detailed descriptions of the videos, we effectively complete the final verification step at a significantly lower cost. After verifying, there are $\mathbf{1,091,712}$ remaining clips. In Fig. 5 (c), we present the statistical distribution of clips duration.
Video quality assessment. To further support research in text-to-video generation, we evaluate the quality of each clip in VideoUFO. Specifically, we adopt six different video quality assessment metrics from VBench [7], which automatically assess videos and align well with human perception. The evaluation dimensions are: (1) subject consistency: assesses whether the main subject’s identity and appearance remain consistent throughout the video; (2) background consistency: evaluates the temporal stability and uniformity of the video’s background; (3) motion smoothness: measures the continuity and fluidity of movements within the video; (4) dynamic degree: assesses the level of activity and variation in motion present in the video; (5) aesthetic quality: judges the visual appeal and attractiveness of the video; and (6) imaging quality: examines the clarity, brightness, and color accuracy of the video. The statistical distributions of scores assessed by these six metrics are shown in the Supplementary (Fig. 9).
Extension. Future researchers can easily extend our VideoUFO in three aspects: (1) Scaling up. Although our VideoUFO has already reached a million-scale level, future researchers may still want to scale it up to 10 million or even 100 million. They can easily achieve this by leveraging our extracted topics to search more videos on platforms such as YouTube and TikTok. (2) New focusing. When we curate VideoUFO, the only available text-to-video prompt dataset is VidProM [35], which is used to analyze user focus and preferences. In the future, user focus may change, and other text-to-video prompt datasets may emerge for analyzing it. Future researchers can easily use our topic extraction pipeline to study these new focuses. (3) Image-to-video. Our current focus is text-to-video, which is the most common approach in the video generation community. Meanwhile, we also notice that image-to-video is gaining popularity. Future researchers can use text and image prompts, such as those in TIP-I2V [34], to analyze the focus of image-to-video users and collect corresponding datasets.
Table 1: The comparison between the proposed VideoUFO with other recent video datasets based on their fundamental attributes. Unlike previous collections, our VideoUFO is derived directly from real user-focused topics and also offers a more flexible license, while remaining comparable to these datasets in other key aspects. “##\##” and “$\ \bar{}$ ” are abbreviations for “numbers” and “average”, respectively.
  Dataset
Venue
Caption
##\##Vid. $\bar{}$ $\bar{}$ Resolution
Domain
##\##Topic
License
WebVid-10M [3] ICCV21 Download 10M 18.0s 14.2 $<$360p Open 1,000 Retracted
HD-VILA-100M [41] CVPR22 Recognition 103M 13.4s 32.5 720p Open 648 R-UDA
InternVid [5] ICLR24 Generated 234M 11.7s 17.6 720p Open 1,051 Apache 2.0
Panda-70M [5] CVPR24 Generated 70M 8.5s 13.2 720p Open 719 R-UDA
LVD-2M [40] NeurIPS24 Generated 2M 20.2s 88.7 Diverse Open 814 R-UDA
MiraData [9] NeurIPS24 Generated 0.33M 72.1s 318.0 720p Open 639 GPL 3.0
Koala-36M [33] Arxiv24 Generated 36M 17.2s 202.1 720p Open 724 R-UDA
VidGen-1M [29] Arxiv24 Generated 1M 10.6s 89.3 720p Open 835 R-UDA
OpenVid-1M [20] ICLR25 Generated 1M 7.2s 127.3 Diverse Open 671 R-UDA
VideoUFO Ours Generated 1M 12.6s 155.5 720p Users $\mathbf{1,291}$ CC BY
4 Comparison with Other Video Datasets
This section compares the proposed VideoUFO with other recent video datasets in terms of fundamental attributes and topics coverage. These differences underscore the necessity of introducing VideoUFO for text-to-video generation.
4.1 Fundamental Attributes
From the Table 1, we draw three conclusions:
  • VideoUFO is collected in line with real text-to-video users’ focus or preferences. In contrast to VideoUFO, other datasets collect videos from the open domain, which may not cover the topics users focus on, and text-to-video models trained on them may fail to meet users’ needs.
  • VideoUFO is released under a more flexible license and introduces new data. Specifically, we search for videos on YouTube by ourselves and only select those with a Creative Commons license. In contrast, most recent datasets (including Panda-70M [5], LVD-2M [40], Koala-36M [33], VidGen-1M [29], and OpenVid-1M [20]) directly source data from HD-VILA-100M [41]. As a result: (1) they do not contribute new data but rather introduce a new data processing pipeline; (2) they must adhere to the same license as HD-VILA-100M, i.e., the Research Use of Data Agreement (R-UDA), which restricts commercial use. In addition, one of the most widely used datasets, WebVid-10M [3], has been retracted due to potential copyright infringement.
  • VideoUFO inherits best practices from other curated datasets. We observe that recent datasets (1) feature detailed captions generated by multimodal large language models (MLLMs), (2) contain millions of video clips, and (3) are high-resolution (i.e., 720p). Therefore, to provide a large-scale and high-quality resource, our VideoUFO inherits these features.
4.2 Topics Coverage
In this section, we analyze the differences in topics coverage between the proposed VideoUFO and existing datasets.
Calculation process. We calculate the number of user-focused topics covered by the existing datasets as follows:
  • Extract topics of recent video datasets. We repeat VidProM’s topic extraction process (Step 1 in Section 3) on the video datasets listed in Table 1. The number of topics extracted for each dataset is shown in Table 1 (##\##Topics).
  • Match the extracted topics with users’ focused ones. We observe that the same or similar topic may be described using different words. For example, both cathedral and church refer to a similar topic. Therefore, when comparing user-focused topics with those from each existing dataset, we choose semantic matching rather than a word-to-word approach. Specifically, (1) we first use $\mathtt{SentenceTransformers}$ [26] to embed the two lists of topics; and (2) then for each user-focused topic, if there exists a topic in the existing dataset with a cosine similarity greater than $0.6$, we consider that user-focused topic to be covered by the existing dataset. Note that the threshold of $0.6$ is an empirical value, as we observe that most similar topics exhibit a cosine similarity greater than $0.6$.
Figure 6: The number of user-focused topics covered by recent video datasets. None successfully includes all user-focused topics.
Figure 7: The calculation process of BenchUFO. It is designed to evaluate whether a text-to-video model can effectively generate videos that contain user-focused topics. It comprises 791 concrete noun topics, each paired with $10$ real-world user-provided prompts.
Observations. The experimental results in Fig. 6 shows:
  • None of the existing datasets cover all user-focused topics in VideoUFO. Specifically, the most comprehensive dataset, InternVid [38], covers $821$ topics, accounting for $63.6\%$ of all user-focused topics. Furthermore, if we combine all existing datasets in Table 1 into the “largest” one, the coverage of user-focused topics will increase to $79.5\%$. However, this will lead to a higher computational burden when training text-to-video models.
  • Interestingly, a dataset may cover more user-focused topics than the actual number of topics it contains. For instance, OpenVid-1M [20] contains $671$ topics but covers $703$ user-focused ones. This is reasonable because two topics in our dataset, VideoUFO, may be similar and thus mapped to a single topic in OpenVid-1M [20]. Nevertheless, this does not compromise the quality of our VideoUFO since it only leads to some relative duplicated topics, while the videos in VideoUFO still accurately reflect real users’ focus and interests.
Limitation. The above topic analysis is based directly on the captions/descriptions provided by each recent video dataset. This is because downloading and re-captioning all these videos is prohibitively expensive in terms of network bandwidth and computation. The quality of these captions/descriptions may affect the number of extracted topics, as some may not be informative. To address this limitation, in the next section, we demonstrate that text-to-video models trained on these datasets fail to generate satisfactory videos for some topics.
5 VideoUFO Benefits Video Generation
5.1 Benchmark
This section details the proposed benchmark, BenchUFO, for evaluating text-to-video models’ performance on user-focused topics. We first introduce the design of the benchmark and then explain why it is reasonable.
Construction. As shown in Fig. 7, the calculation process of our BenchUFO includes: (1) Selecting prompts. First, we select 791 concrete nouns from $1,291$ user-focused topics. (Please refer to the next section for the rationale behind selecting concrete nouns.) Then, for each chosen topic, we randomly select $10$ text prompts from VidProM [35]. All these prompts constitute the prompt set for our benchmark. (2) Generating videos. For each prompt, we use a text-to-video model to generate a corresponding video. (3) Describing videos. For each video, we use a multimodal large language model (here, we choose $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ [32]) to understand and describe it. The instruction prompt is provided in the Supplementary (Section 10). (4) Calculating similarity. We use a sentence embedding model (in this case, $\mathtt{SentenceTransformers}$ [26]) to encode both the input prompt and the output description, and then compute the cosine similarity between them. We calculate the average similarity across $10$ prompts for each topic. A higher similarity score indicates better performance on that specific topic. Our benchmark computes and considers the $10$/$50$ worst-performing and $10$/$50$ best-performing topics.
Table 2: The performance of both publicly available text-to-video models and our trained models on the proposed BenchUFO. The publicly available models are trained on various datasets, including both public and private ones. MVDiT [20] is trained on VidGen [29], OpenVid [20], and VideoUFO, respectively. “Low/Top $N$” denotes the average score of the worst/best-performing $N$ topics.
  Models
Low $10$
Low $50$
Top $50$
Top $10$
Mira [10] $0.236$ $0.282$ $0.508$ $0.550$
Show-1 [44] $0.266$ $0.303$ $0.524$ $0.564$
LTX-Video [15] $0.268$ $0.310$ $0.532$ $0.574$
Open-Sora-Plan [16] $0.314$ $0.361$ $0.559$ $0.598$
TF-T2V [36] $0.316$ $0.359$ $0.560$ $0.595$
Mochi-1 [30] $0.323$ $0.367$ $0.580$ $0.616$
HiGen [24] $0.352$ $0.394$ $0.589$ $0.625$
Open-Sora [45] $0.363$ $0.409$ $0.601$ $0.639$
Pika [22] $0.365$ $0.404$ $0.583$ $0.619$
RepVideo [28] $0.368$ $0.402$ $0.589$ $0.619$
T2V-Zero [11] $0.375$ $0.410$ $0.586$ $0.616$
CogVideoX [42] $0.383$ $0.419$ $0.601$ $0.629$
Latte-1 [18] $0.384$ $0.421$ $0.592$ $0.636$
HunyuanVideo [12] $0.388$ $0.427$ $0.612$ $0.645$
LaVie [37] $0.399$ $0.426$ $0.595$ $0.632$
Pyramidal [8] $0.400$ $0.433$ $0.606$ $0.647$
MVDiT-VidGen [29] $0.382$ $0.426$ $0.594$ $0.626$
MVDiT-OpenVid [20] $0.401$ $0.437$ $0.609$ $0.645$
MVDiT-VideoUFO $\mathbf{0.442}$ $\mathbf{0.465}$ $\mathbf{0.619}$ $\mathbf{0.651}$
Justification. This section provides justification for the proposed BenchUFO from three perspectives:
  • Why choose concrete nouns? This is due to the inconsistency between the input prompt and the output description when using abstract nouns. For instance, for the abstract noun topic “freedom” and the prompt “In a freedom world, ideal and love both exist”, a model may generate videos containing “dove”, “star”, or “heart”, and the video understanding model will summarize them as “a dove is playing…”, “a star is shining…”, and “a heart is beating…”. However, the embeddings of these descriptions are not expected to be close to those of the given prompt. This might create a false impression that the model cannot effectively generate content for that topic.
  • Why use video understanding model reasonable? This is because the video understanding model has advanced significantly with the advent of large language models, and it is reasonable to assume that these models have been exposed to images/videos containing user-focused topics. Here, we choose $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ [32] as the video understanding model. Trained on $800$ billion tokens of visual-related data and $600$ billion tokens of text data, it is expected to accurately comprehend these user-focused topics and faithfully summarize a video.
  • Why not use established benchmarks? This is because the existing benchmarks fail to reflect real-world scenarios.. Specifically, we note that there exists a well-established benchmark, VBench [7], whose “object class” and “multiple objects” dimensions are similar to our BenchUFO. However, the prompts in VBench have two main drawbacks: (1) They cover only a limited number of topics/objects – specifically, just $79$. (2) They focus exclusively on common topics/objects (e.g., “bicycle”, “car”, and “airplane”), while many topics that users care about, such as “forest”, “sunset”, and “beach”, are missing.
Figure 8: Visual comparisons between our approach (MVDiT-VideoUFO) and other text-to-video models. The model trained on VideoUFO outperforms the alternatives in generating user-focused topics. In the prompts, concepts in green indicate those successfully generated solely by our method, whereas those in yellow denote concepts successfully generated by both our model and the competing models. Here we represent each video with a single frame, and all generated videos are provided in the supplementary materials.
5.2 Observations
In this section, we evaluate current text-to-video models on this new benchmark and show that, with the help of VideoUFO, we achieve state-of-the-art performance. The quantitative and qualitative results are shown in Table 2 and Fig. 8, respectively. We observe that:
  • Current text-to-video models do not consistently perform well across all user-focused topics. Specifically, there is a score difference ranging from $0.233$ to $0.314$ between the top-$10$ and low-$10$ topics. These models may not effectively understand topics such as “giant squid”, “animal cell”, “Van Gogh”, and “ancient Egyptian” due to insufficient training on such videos.
  • Current text-to-video models show a certain degree of consistency in their best-performing topics. We discover that most text-to-video models excel at generating videos on animal-related topics, such as ‘seagull’, ‘panda’, ‘dolphin’, ‘camel’, and ‘owl’. We infer that this is partly due to a bias towards animals in current video datasets.
  • The proposed VideoUFO helps reduce the gap between the worst-performing and best-performing topics. To demonstrate the effectiveness of the proposed VideoUFO, we train an MVDiT [20] solely on VideoUFO. We find that the trained model achieves the highest low-$10$ scores (a $+4.2\%$ improvement compared to the current state-of-the-art) while maintaining performance on the top-$10$ topics. Visually, the trained model successfully generates videos on topics that other models previously could not.
  • The proposed VideoUFO outperforms other similar-scale datasets, such as OpenVid [20] and VidGen [29]. To demonstrate that the improvement originates from our VideoUFO rather than from the MVDiT architecture [20], we replace our VideoUFO with OpenVid [20] and VidGen [29]. We find that when using the popular OpenVid [20], the low-$10$ scores are similar to previous state-of-the-art models ($0.401$ vs. $0.400$). Furthermore, when training on VidGen [29], the performance is even lower.
5.3 Future works
Due to limited computational resources, when validating the effectiveness of our dataset, we currently follow the video generation method from the academic community introduced by [20], which already requires $32$ A100 GPUs. However, the quality of videos generated by this baseline remains limited. In the future, industry researchers could incorporate our dataset into their large-scale foundational model training pipelines to further explore its potential.
6 Ablation Studies on Curation of VideoUFO
This section performs ablation studies to investigate key components in curating VideoUFO, i.e., examining (1) the impact of the number of clusters in user-focused topic analysis and (2) the methods for clip verification. For more details, please refer to Supplementary (Section 11).
7 Conclusion
This paper presents a newly collected million-scale dataset for text-to-video generation with a focus on user needs. Beyond this, our dataset exhibits minimal overlap with existing datasets and is released under a more permissive license (i.e., CC BY). We first describe the curation process, which involves analyzing user-focused topics, searching for these topics, segmenting the resulting videos, and applying various post-processing techniques. Then, we highlight the differences between our dataset and existing ones in terms of fundamental attributes and topics coverage. Finally, we build a new benchmark to evaluate text-to-video models on user-focused topics and demonstrate that our dataset helps improve model performance in this regard. We encourage both the research community and industry to use our dataset to further advance the field of text-to-video generation.
Acknowledgment
We sincerely thank OpenAI for their support through the Researcher Access Program. Without their generous contribution, this work would not have been possible.
Supplementary Material
Figure 9: The statistical distributions of scores assessed by these six video quality metrics, which have been aligned with human perception. Researchers can use these scores to filter videos according to their specific training needs.
8 The Prompt for Concluding Topics
Prompt: Could you describe the topic in the following short sentences using only 1-2 words? Please return only the topic (1-2 word) as a singular noun or in the base form of the verb.
9 The Distributions of Video Quality Score
Fig. 9 shows the statistical distributions of scores evaluated in terms of subject consistency, background consistency, motion smoothness, dynamic degree, aesthetic quality, and imaging quality.
10 The Prompt for Describing Videos
[…]
11 Ablation Studies on Curation of VideoUFO
This section performs ablation studies to investigate key components in curating VideoUFO, i.e., examining (1) the impact of the number of clusters used in user-focused topic analysis and (2) the methods used for clip verification.
The number of clusters. As shown in Fig. 10, we vary the number of pre-set clusters from $100$ to $3,000$ when analyzing the focused topics of text-to-video users. We observe that initially, as the number of pre-set clusters increases, the number of resulting topics also increases; however, after reaching approximately $2,000$ clusters, this increase stops and begins to fluctuate. Therefore, we finally choose $2,000$ as the number of pre-set clusters. This is reasonable because the number of user-focused topics is limited, and an excessive number of pre-set clusters would eventually merge together. Beyond this, we also experiment with DBSCAN and HDBSCAN [19] to automatically determine the number of clusters; however, we do not obtain reasonable results.
The method for clip verification. We randomly select $1,000$ video clips to conduct ablation studies on the methods used to verify whether a clip contains a specific topic. Feeding the entire video into GPT-4o (denoted as “Direct GPT-4o”) is considered the most accurate approach, and we compare the overlap of other cost-effective methods with it. As shown in Table 3, we observe that: (1) Feeding videos directly into GPT-4o is very expensive – processing $1,000$ clips costs \$$\$162.23$. (2) Using detailed captions instead of video clips reduces the API cost by approximately 5,400×5,400$\times$5 , 400 ×, while maintaining about $80\%$ prediction overlap. (3) Although using brief captions is even cheaper, the overlap drops by $14.9\%$, which is reasonable since a brief caption cannot capture all the information in a clip. (4) Interestingly, we find that GPT-4o-mini achieves slightly better performance than GPT-4o. In conclusion, for verifying clips, we choose detailed captions with GPT-4o-mini.
Figure 10: The relationship between pre-set number of clusters for K-means and final number of user-focused topics.
Table 3: The methods for clip verification. We use $1,000$ clips for this ablation study. “Direct GPT-4o” refers to processing videos directly with GPT-4o, and “Overlap” measures the extent to which the predictions of other methods match those of “Direct GPT-4o”.
  Method ##\## Verified Overlap API Cost
Direct GPT-4o $507$- \$$\$162.23$
Brief GPT-4o $174$ $0.621$$\$0.04$
GPT-4o-mini $276$ $0.657$$\$0.01$
GPT-4o $299$ $0.770$$\$0.46$
Detail GPT-4o-mini $393$ $0.806$$\$0.03$
12 The List of Full Topics
  dance horror alien cat
music fashion walk forest
cyberpunk portrait flight battle
drive nature rain sunset
dog celebration romance beach
mythology art explosion conflict
eat zombie run landscape
moon robot space village
dragon cityscape lion robotics
rabbit race desert night
christmas bird cook culture
paint girl visualization beauty
animal play fire magic
fish war jungle game
car architecture boy astronomy
coffee mountain weather warrior
snow logo wave monkey
swim friendship haunt cloud
ufo smoke garden gothic
woman cartoon chase
cinematography
flower conversation ghost angel
elephant psychedelia superheroes future
  vehicle scenery dinosaur humor
celebrity underwater sunrise history
combat ocean ritual apocalypse
emotion meditation demon bear
wildlife transportation cyborg warfare
mystery light fairy monster
motorcycle royalty tree tiger
wolf sleep family mermaid
batman futurism comedy tourism
travel panda lake water
egypt train drone interaction
astronaut horse child soccer
destruction design samurai
chinese culture
frog smile religion sing
advertise mar sky winter
butterfly fox cosmos wind
hair box aviation growth
pirate festival steampunk guitar
vampire censorship medicine mushroom
halloween couple love spaceship
shark soldier exercise eye
read river fruit martial arts
luxury snake kitten urbanization
mouse elon musk santa claus castle
spacecraft atmosphere basketball temple
indian culture
loneliness man storm
squirrel nightlife ant office
action spirituality noir star wars
iron man timelapse meme metropolis
hinduism imagery chicken survival
artificial intelligence
fall decoration performance
whisker football pig santa
princess city sadness paranormal
crime insect automobile arte
duck flag nostalgia accident
neon study waterfall tema
shop turtle viking supernatural
wed jesus concert character
cow witch interior journey
ninja dreamscape owl rap
goddess literature news boat
jump clown nighttime psychedelics
disney tranquility cannabis medieval
volcano firearm pikachu dj
storytelling filmmaking spiderman ship
monk expression
swordsmanship
rainforest
barbie skateboard piano market
marine life military model surf
ice cream school television infant
firework neuroscience creature vlogging
sculpture the matrix experiment time travel
party gold bible park
camera movement
meteor octopus whale
cycle space battle eagle autumn
din spider-man crash finance
door skyscraper launch tea
escape face
food combination
fight
arab culture cemetery room ai
  time police spider bloom
crow conversion wealth doll
nightclub phone botany apple
night sky sport trump balloon
pizza earth
cryptocurrency
crystal
parody banana rotation environment
black hole heart dolphin planet
death
haunted mansion
superman knight
dragonfly giant russia krishna
cake bunny countryside statue
politics cowboy origami gather
childhood penguin drift act
new york reflection perseverance star
dystopia cricket
horseback riding
jellyfish
routine classroom sheep rise
money climb witchcraft deity
chaos laughter fractal evolution
snowfall clothe treasure flying car
cosmology skeleton king pyramid
street glass fear candy
dream ancient rome color agriculture
prayer book meet disaster
island
virtual reality
sea creature tokyo
parent teddy bear investigation food
traffic bicycle farmer jewelry
gangster violence commercial encounter
advertisement bollywood youth teenager
symbolism pixar aesthetic solar
deer wizardry victory neighborhood
mutation crustacean library puppy
religious imagery
portal cafe crossover
china chess elf parrot
construction gameplay hell pollution
ancient greece
greet bounce ambiance
gallop sound toy ballet
robot combat attire videogame beer
fitness costume heaven islam
iceberg paris werewolf hospital
footage drum satan cat behavior
sun galaxy blink rock
mysticism simulation civilization mask
bar trade archaeology divinity
relationship farm graffiti dentistry
lionel messi employment protest assassination
toilet
slavic mythology
youtube wine
education homelessness shadow resilience
mansion bakery vegetable gamble
sail ceremony unicorn speed
marvel gorilla africa visual effects
skyline humanoid godzilla clock
clean write chocolate realism
sci-fi egg tropical thriller
yoga cuisine dark bodybuilding
mirror hack determination thunderstorm
rave turkey bridge elderly
abandonment puppetry goat appearance
minecraft render rainbow be
sitcom reunion bedroom house
  hulk anger taylor swift hybridization
hunt
communication
happiness meadow
background field skull dinner
lightning tennis candle intelligence
swimming pool
energy disco relaxation
gift business bus illness
seaside program alcohol disapproval
imprisonment kitchen gesture cry
punk hand nutrition social media
record floral life soviet
truck aerial draw shiva
compassion digital ski bubble
italy retro computer cinema
wrestle fog tribalism hallway
crocodile restaurant lego cybersecurity
giraffe
cristiano ronaldo
speech camp
sand introspection
character design
peacock
textile raccoon joy cuteness
victoriana spaghetti church manufacture
blonde living room skydive fatherhood
heroism moonlight satire helicopter
picnic bikini safety extraterrestrial
leaf motivation telephone hotel
bee spongebob occult nuclear
marine astrology teach photoshoot
tornado hug display asteroid
cherry blossoms
prehistory health robbery
aurora
music production
hip-hop darkness
tattoo brazil breakdancing darth vader
goku flood suspense geometry
digital art science student circus
emergency ganesha joker cheese
behavior stargaze theater quantum
waste physic delivery folklore
pokemon shoe holography card
mexico coral reefs machinery villain
sneaker tesla dive prophet
woodworking palace
anthropomorphism
surveillance
coin burger blue muppets
korea diplomacy calendar antique
einstein campfire presentation flame
superpower bottle grandmother bedtime
message collaboration interview drug
pixel art jazz body electricity
anatomy rescue terrorism sea monster
primate donkey brand skate
mentorship beverage dog eating lizard
burn cabin wonderland event
makeup milk archery connectivity
rebellion splash synthesizer pumpkin
eiffel tower surprise kiss contemplation
urban buddhism mathematics bully
map mecha
nuclear warfare
chemistry
wisdom tunnel lo-fi buddha
science fiction
procrastination
print farewell
genetics
mercedes-benz
indonesia fable
foot diamond secrecy duel
dynasty cell philosophy economics
  bmw metalworking strawberry biblical
golf hope abstraction weightlift
camel gratitude colonialism
bioluminescence
corgi crowd breakfast violin
cliff bat exhibition signage
dubai
fluid dynamics
morocco bag
diversity selfie empowerment worm
climate afterlife
amusement park
malfunction
geology microbiology
head movement
imran khan
cybernetics
public speaking
stock market purple
shrek intimacy sibling maya
android pond firefighting pregnancy
tsunami leadership iran mcdonald’s
confusion aztec lightsaber hydration
mosque baseball kingdom sphere
soul surgery perspective muslim
maze
mickey mouse
mario
mental health
gender espionage conspiracy muscle
techno security levitation creativity
seascape vehicle alphabet grasshopper
repair parade style tarot
shepherd migration cooky vaporwave
injury season furniture metal
potato firefly wake illusion
silhouette treehouse bath trailer
companionship
skincare terminator super mario
pakistan kungfu argument psychology
bullfighting perfume melt data
oil watermelon heal antarctica
empire earthquake university phoenix
grief himalaya entertainment wall
real estate putin discovery redhead
jugando afrofuturism soda hanuman
slavery cottage kaiju tyrannosaurus
sunshine businessman hong kong glitch
laptop diwali
street photography
hedgehog
carnival orange napoleon wash
community awaken
mortal kombat
skin
law gladiator blood justice
noah’s ark anxiety hamster spain
heist tolkien stadium bigfoot
tank charity fungus mammoth
eminem pope craft cinderella
chair noodle hygiene biomechanics
deception pandemic courtroom pug
fast food sunlight titanic hamburger
campus supercar children’s art rooster
hotdog thailand temperature
naval warfare
expressionism typography luck polar bear
charge type snail swamp
laser ethereal subterranean doraemon
riot confrontation mahabharata predation
school bus zoo ceo mongol
monroe sushi collision ink
lgbtq+ superhero goblin spartan
donut atlantis
web development
minimalism
panther iphone cryptids e-commerce
submarine lip-sync beard vend
  lava stress airport dessert
yacht particle capybara lotus
celtic prince caterpillar nurse
nebula indoors aquarium mujer
lemon
ottoman empire
fireplace warhammer
radio virus cube currency
cup examination infidelity courage
honey transformer unity connection
baldness arrest influencer proposal
lifestyle knife gas station taxi
content
metamorphosis
entrepreneurship
gandhi
sandwich slime suit raven
biology emoji shower spacewomen
dog smiling crucifixion priest nativity
minion emergence taco language
graduation ballistics inspiration alley
cleopatra kidnap hill preach
thor consultation prophecy fan
cub plague dracula motherhood
childbirth kangaroo sonic peru
businesswoman
saturn burnout juice
watch spice dunk aspiration
square umbrella automation puzzle
anticipation hallucination navigation mandala
rama search breathe wish
cannibalism plumb fairytale mine
sloth engine seagull australia
pineapple logistics blindness backroom
impact pottery mindfulness customer
alligator
decision-making
greed hockey
indoor long rollercoaster transition
snack deadpool game show volleyball
laundry cheetah alexander cocktail
existentialism sikhism imagination alchemy
scorpion squid
schwarzenegger
geography
decay massage freedom fatigue
cosplay adam driver tupac hippy
llama interracial chameleon dust
knit gundam
nursery rhyme
curse
nail patriotism handshake sniper
mughals notification profession workspace
augmented reality
pool party morality popcorn
luffy commute
shawn mendes
independence
snoop dogg contrast observation mutant
swan monastery digestion group
abraham interrogation vote seahorse
drought taj mahal comet therapy
niño teletubbies caminar jupiter
positivity ladybug sofa olympics
drink salad homecoming vase
confidence cuba remote work circulation
soup cancer stormtrooper venom
blade runner consciousness addiction gingerbread
sew
astral projection
sri lanka philippine
fry lecture peach forage
retrofuturism pastry harley quinn gravity
blacksmithing lily inquisitor rugby
landmark ramayana
genshin impact
backflip
  zelda urban decay tea ceremony toast
chaplin asmr elvis hypnosis
curtain throw marathon odyssey
persona teleportation theory temptation
kurt cobain rapunzel hatsune miku destiny
division bass space shuttle yin-yang
billie eilish
sleep paralysis
horizon garfield
biblical-figures
mia khalifa dua lipa
unreal engine
half-life dieselpunk chuck norris eggplant
tower
deepika padukone
puppeteering limitation
seinfeld oracle garvey