### v2

# VideoUFO: A Million-Scale User-Focused Dataset for Text-to-Video Generation

Abstract
Text-to-video generative models convert textual prompts into dynamic visual content, offering wide-ranging applications in film production, gaming, and education. However, their real-world performance often falls short of user expectations. One key reason is that these models have not been trained on videos related to some topics users want to create. In this paper, we propose VideoUFO, the first Video dataset specifically curated to align with Users’ FOcus in real-world scenarios. Beyond this, our VideoUFO also features: (1) minimal ($0.29\%$) overlap with existing video datasets, and (2) videos searched exclusively via YouTube’s official API under the Creative Commons license. These two attributes provide future researchers with greater freedom to broaden their training sources. The VideoUFO comprises over $1.09$ million video clips, each paired with both a brief and a detailed caption (description). Specifically, through clustering, we first identify $1,291$ user-focused topics from the million-scale real text-to-video prompt dataset, VidProM. Then, we use these topics to retrieve videos from YouTube, split the retrieved videos into clips, and generate both brief and detailed captions for each clip. After verifying the clips with specified topics, we are left with about $1.09$ million video clips. Our experiments reveal that (1) current $16$ text-to-video models do not achieve consistent performance across all user-focused topics; and (2) a simple model trained on VideoUFO outperforms others on worst-performing topics. The dataset and code are publicly available here and here under the CC BY 4.0 License.
Figure 1: VideoUFO is the first dataset curated in alignment with real-world users’ focused topics for text-to-video generation. Specifically, the dataset comprises over $1.09$ million video clips spanning $1,291$ topics. Here, we select the top $20$ most popular topics for illustration. Researchers can use our VideoUFO to train or fine-tune their text-to-video generative models to better meet users’ needs.
1 Introduction
Text-to-video generation aims to convert textual descriptions into dynamic visual content. Its applications are extensive and transformative, covering areas from creative media bender2023coexistence and entertainment yu2024barriers to practical domains such as education garcia2024pre , advertising zheng2024open , and assistance li2024endora .
Figure 2: The glowing firefly: (a) generated by Sora openai2024sora and (b) captured in a real video. The generated firefly is noticeably different from its real-life counterpart and thus unsatisfying. We attribute this primarily to a lack of exposure to such topics.
Despite their popularity and usefulness, current text-to-video models often fail to meet users’ expectations in real-world applications. For example, when we ask Sora openai2024sora to generate a video using the prompt “A firefly is glowing on a grass’s leaf on a serene summer night”, it fails to capture the concept of a glowing firefly while successfully generating grass and a summer night, as shown in Fig. 2 (a). From the data perspective, we infer this is mainly because Sora openai2024sora has not been trained on firefly-related topics, while it has been trained on grass and night. Furthermore, if Sora openai2024sora has seen the video shown in Fig. 2 (b), it will understand what a glowing firefly should look like.
To this end, this paper presents VideoUFO, the first Video dataset curated specifically based on real-world Users’ FOcus in text-to-video generation. Such a dataset can help improve the alignment between text-to-video models and actual user needs. Specifically, we focus on: (1) curating VideoUFO by analyzing real users’ interests and scraping relevant videos; (2) comparing VideoUFO with other datasets to highlight differences; and (3) showing how VideoUFO benefits video generation.
The first dataset that aligns with real-world users’ focus in text-to-video generation. Our key idea is to analyze user-focused topics from user-provided prompts and then search for videos related to these topics. As shown in Fig. 1, the resulted VideoUFO comprises more than $\mathbf{1.09}$ million video clips spanning $\mathbf{1,291}$ user-focused topics. Specifically, we initiate by analyzing user-focused topics, which involves embedding all $1.67$ million user-provided prompts from VidProM wang2024vidprom , clustering these embeddings using K-means, and generating a topic for each cluster with GPT-4o openai2024hello , followed by combining similar topics. After obtaining these topics, we (1) search for these topics on YouTube, (2) segment the fetched videos into multiple semantically consistent short clips, (3) generate both brief and detailed captions for each clip, (4) filter out clips that do not contain the specific topics, and (5) assign each clip video quality scores that align with human perception. Note that although our current VideoUFO comprises about one million videos, it can be easily scaled up to ten million or more by sourcing additional videos for each topic. The VideoUFO can also be easily extended to the image-to-video domain by using the text and image prompts in TIP-I2V wang2024tip .
Differences between VideoUFO and other recent video datasets. Recently, several video datasets have been released, including OpenVid-1M nan2024openvid , HD-VILA-100M xue2022advancing , InternVid wanginternvid , Koala-36M wang2024koala , LVD-2M xiong2024lvd , MiraData ju2024miradata , Panda-70M chen2024panda , VidGen-1M tan2024vidgen , and WebVid-10M bain2021frozen . While inheriting their excellent attributes – such as large scale, accurate captions, and high resolution – our VideoUFO explores some novel directions: (1) Guided by real user focus. Specifically, whereas recent datasets are typically gathered from open-domain sources, we concentrate on topics that are focused by text-to-video users. This feature enables text-to-video models trained on our dataset to better cater to users, while avoiding unnecessary expansion of the dataset and wasting resources. (2) Introducing new data. Although these recent papers claim to contribute new datasets, most of them primarily introduce a new data pipeline – that is, they reprocess HD-VILA-100M xue2022advancing and are subsets of it. While their contributions are useful and meaningful, in theory, a generative model that has already been fully fitted on HD-VILA-100M xue2022advancing would not gain new video knowledge from them. In contrast, we collect new data from YouTube, with only $0.29\%$ of the videos overlapping with existing datasets. (3) Data compliance. When curating VideoUFO, we retrieve videos using YouTube’s official API and select only those with a Creative Commons license. In contrast, most recent datasets do not explicitly address the regulatory compliance of their data collection process. This feature grants researchers greater flexibility in using our data.
Benchmarking current text-to-video models on user-focused topics and demonstrating the effectiveness of our VideoUFO. We present a new benchmark to quantify the performance of text-to-video models on user-focused topics. Specifically, our process involves: (1) selecting $10$ user-provided prompts per topic; (2) generating a video for each prompt; (3) using a multimodal large language model to describe each video; and (4) calculating the similarity between the generated descriptions and the original prompts. For each topic, we calculate the average similarity between the $10$ prompts and their corresponding descriptions. By sorting these averages, we can identify the worst-performing and best-performing topics for each model. Using this benchmark, we evaluate the performance of current text-to-video models and a newly trained text-to-video model on our VideoUFO. Experimental results indicate that (1) current $16$ text-to-video models have some poor-performing topics, and (2) our model achieves the highest similarities on worst-performing topics while maintaining performance on the best-performing ones.
In conclusion, our key contributions are as follows:
  1. 1.
    We present VideoUFO, the first video dataset curated based on the focus of real text-to-video users. This dataset comprises over $1.09$ million clips spanning $1,291$ user-focused topics.
  2. 2.
    We compare VideoUFO with recent video datasets, highlighting their differences in both fundamental attributes and topics coverage, thereby emphasizing the necessity of our dataset. We also follow best practices in their curation processes to ensure the quality of our dataset.
  3. 3.
    We evaluate current text-to-video models on user-focused topics and observe that a simple model trained on our VideoUFO outperforms competing models on worst-performing topics.
2 Related Works
Text-to-Video Generation. The introduction of Sora openai2024sora has sparked significant research interest in text-to-video generation. Commercial models such as Movie Gen polyak2024moviegen , Veo veo2024 , Kling kuaishou_kling , PixelDance bytedance2024pixeldance_seaweed , and Seaweed bytedance2024pixeldance_seaweed demonstrate strong performance and are increasingly being integrated into various industries. Meanwhile, open-source models like HunyuanVideo kong2024hunyuanvideo , LTX-Video ltxvideo , CogVideoX yang2024cogvideox , Mochi-1 genmo2024mochi , and Pyramidal jin2024pyramidal empower researchers to experiment, customize, and enhance existing frameworks. Although they are powerful, their training data determines the upper limit of generated videos’ quality. This paper proposes a dataset, VideoUFO, which has the potential to help these models better cater to users’ preferences in real-world applications.
Figure 3: Each data point in our VideoUFO includes a video clip, an ID, a topic, start and end times, a brief caption, and a detailed caption. Beyond that, we evaluate each clip with six different video quality scores from VBench huang2024vbench .
Text-Video Datasets. A text-video dataset consists of video clips paired with corresponding textual descriptions or captions. Many text-video datasets, such as OpenVid-1M nan2024openvid , HD-VILA-100M xue2022advancing , InternVid wanginternvid , Koala-36M wang2024koala , LVD-2M xiong2024lvd , MiraData ju2024miradata , Panda-70M chen2024panda , VidGen-1M tan2024vidgen , and WebVid-10M bain2021frozen , have been proposed to advance multimodal understanding and generation. However, these datasets are not designed to align with the preferences of text-to-video users, creating a gap between training data topics and real-world applications. In contrast, our dataset VideoUFO is collected from a user-focus perspective. To better cater users, future researchers can fine-tune text-to-video models with the VideoUFO or integrate it with existing datasets to develop new models.
Preference Alignment in Diffusion Models. While diffusion models have demonstrated impressive capabilities, there is an increasing demand to ensure their outputs align with human preferences. Direct Preference Optimization (DPO) rafailov2023direct , originally proposed to align large language models (LLMs) with human preferences, can also be effectively applied to diffusion models. For example, DiffusionDPO wallace2024diffusion fine-tunes text-to-image models using human comparison data to improve image generation quality. Additionally, VideoDPO liu2024videodpo introduces a comprehensive preference scoring system that evaluates both visual quality and semantic alignment to enhance text-to-video generation. Unlike other approaches that align text-to-image or text-to-video models with attributes such as aesthetics, motion, consistency, and visual appeal, our goal is to improve the performance of text-to-video models on real-world users’ focused topics.
3 Curating VideoUFO
Fig. 3 illustrates a data point from our million-scale VideoUFO, which comprises: (1) a video clip along with its ID; (2) the topic of the clip; (3) the start and end times within its original video; (4) the corresponding brief and detailed captions; and (5) the quality of the clip. The following provides an explanation of how we curate the VideoUFO.
[…]
Analyzing text-to-video users’ focused topics. We aim to analyze the topics that real users focus on or prefer when generating videos from texts. Here, we use VidProM wang2024vidprom as it is the only publicly available, million-scale text-to-video prompt dataset written by real users. First, we embed all $1.67$ million prompts from VidProM into $384$-dimensional vectors using $\mathtt{SentenceTransformers}$ reimers2019sentencebert . Next, we cluster these vectors with K-means. Note that here we pre-set the number of clusters to a relatively large value, i.e., $2,000$, and merge similar clusters in the next step. Finally, for each cluster, we ask GPT-4o openai2024hello to conclude a topic ($1\sim 2$ words). The prompt is shown in the Appendix (Section A). After merging singular and plural forms, restoring verbs to their base forms, removing duplicates, and manually verifying the topics – removing overly broad ones such as ‘animation’, ‘scene’, ‘movement’, and ‘film’ – we finally obtained $\mathbf{1,291}$ topics. The semantic distribution of these topics is shown in Fig. 4.
Collecting videos from YouTube. For each topic obtained, we use the official YouTube API to search for videos. Specifically, for each topic, we search for approximately $500$ videos by relevance, with each video meeting the following criteria: (1) it is shorter than $4$ minutes, (2) it has a resolution of $720$p or higher, and (3) it is licensed under Creative Commons. These requirements ensure that the collected videos are suitable and freely usable for training video generation models. In the end, we obtain $586,490$ videos from YouTube. We compare the YouTube IDs of these videos with those in existing datasets, including OpenVid-1M nan2024openvid , HD-VILA-100M xue2022advancing , InternVid wanginternvid , Koala-36M wang2024koala , LVD-2M xiong2024lvd , MiraData ju2024miradata , Panda-70M chen2024panda , VidGen-1M tan2024vidgen , and WebVid-10M bain2021frozen , and find that only $1,675$ IDs ($0.29\%$) are already present in these datasets. This suggests that our VideoUFO introduces novel information or knowledge, which can serve to expand the range of existing training sources.
Splitting videos and generating captions. After obtaining the videos, we segment them into multiple semantics-consistent short clips following the steps in curating Panda-70M chen2024panda . The process includes shot boundary detection, stitching, and video splitting, producing a total of $\mathbf{3,181,873}$ clips. To facilitate the training of text-to-video models on our VideoUFO, we generate both brief and detailed captions for each video clip. For the brief captions, we utilize the video captioning model provided by Panda-70M chen2024panda . For the detailed captions, we adopt the pipeline used in $\mathtt{Open}$-$\mathtt{Sora}$-$\mathtt{Plan}$ v1.3.0 lin2024open , which employs $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ Qwen2VL for video annotation. In the Appendix (Fig. 10 (a) and (b)), we present the statistical distribution of the lengths of brief and detailed captions, respectively.
Verifying clips. We notice that, although the videos are searched by topics, not every clip contains our intended topics. A straightforward solution is to use GPT-4o openai2024hello to verify whether a given clip corresponds to a specific topic. However, while effective, this approach would be prohibitively expensive for verifying more than $3$ million video clips. As an alternative, we use the detailed caption (instead of the original clip) and feed it into GPT-4o mini (instead of GPT-4o) to verify whether it belongs to a specific topic. Since (1) GPT-4o mini and GPT-4o have similar capabilities in basic language understanding, and (2) the video understanding model provides detailed descriptions of the videos, we effectively complete the final verification step at a significantly lower cost. After verifying, there are $\mathbf{1,091,712}$ remaining clips. In the Appendix (Fig. 10 (c)), we present the statistical distribution of clips duration.
Table 1: The comparison between the VideoUFO with recent video datasets based on their fundamental attributes. Unlike previous collections, our VideoUFO is derived directly from real user-focused topics and also offers a more flexible license, while remaining comparable to these datasets in other aspects. “##\##” and “$\ \bar{}$” are abbreviations for “numbers” and “average”, respectively.
  Dataset ##\##Vid. $\bar{}$ $\bar{}$ Resolution
Domain
##\##Topic
License
WebVid-10M bain2021frozen 10M 18.0s 14.2 $<$360p Open 1,000 Retracted
HD-VILA-100M xue2022advancing 103M 13.4s 32.5 720p Open 648 R-UDA
InternVid chen2024panda 234M 11.7s 17.6 720p Open 1,051 Apache 2.0
Panda-70M chen2024panda 70M 8.5s 13.2 720p Open 719 R-UDA
LVD-2M xiong2024lvd 2M 20.2s 88.7 Diverse Open 814 R-UDA
MiraData ju2024miradata 0.33M 72.1s 318.0 720p Open 639 GPL 3.0
Koala-36M wang2024koala 36M 17.2s 202.1 720p Open 724 R-UDA
VidGen-1M tan2024vidgen 1M 10.6s 89.3 720p Open 835 R-UDA
OpenVid-1M nan2024openvid 1M 7.2s 127.3 Diverse Open 671 R-UDA
VideoUFO 1M 12.6s 155.5 720p Users $\mathbf{1,291}$ CC BY
Video quality assessment. To further support research in text-to-video generation, we evaluate the quality of each clip in VideoUFO. Specifically, we adopt six different video quality assessment metrics from VBench huang2024vbench , which automatically assess videos and align well with human perception. The evaluation dimensions are: (1) subject consistency: assesses whether the main subject’s identity and appearance remain consistent throughout the video; (2) background consistency: evaluates the temporal stability and uniformity of the video’s background; (3) motion smoothness: measures the continuity and fluidity of movements within the video; (4) dynamic degree: assesses the level of activity and variation in motion present in the video; (5) aesthetic quality: judges the visual appeal and attractiveness of the video; and (6) imaging quality: examines the clarity, brightness, and color accuracy of the video. The statistical distributions of scores assessed by these six metrics are shown in the Appendix (Fig. 11).
Extension. Future researchers can easily extend our VideoUFO in three aspects: (1) Scaling up. Although our VideoUFO has already reached a million-scale level, future researchers may still want to scale it up to 10 million or even 100 million. They can easily achieve this by leveraging our extracted topics to search more videos on platforms such as YouTube and TikTok. (2) New focusing. When we curate VideoUFO, the only available text-to-video prompt dataset is VidProM wang2024vidprom , which is used to analyze user focus and preferences. In the future, user focus may change, and other text-to-video prompt datasets may emerge for analyzing it. Future researchers can easily use our topic extraction pipeline to study these new focuses. (3) Image-to-video. Our current focus is text-to-video, which is the most common approach in the video generation community. Meanwhile, we also notice that image-to-video is gaining popularity. Future researchers can use text and image prompts, such as those in TIP-I2V wang2024tip , to analyze the focus of image-to-video users and collect corresponding datasets.
4 Comparison with Other Video Datasets
This section compares the proposed VideoUFO with other recent video datasets in terms of fundamental attributes and topics coverage. These differences underscore the necessity of introducing VideoUFO for text-to-video generation.
4.1 Fundamental Attributes
From the Table 1, we draw three conclusions:
  • VideoUFO is collected in line with real text-to-video users’ focus or preferences. In contrast to VideoUFO, other datasets collect videos from the open domain, which may not cover the topics users focus on, and text-to-video models trained on them may fail to meet users’ needs.
  • VideoUFO is released under a more flexible license and introduces new data. Specifically, we search for videos on YouTube by ourselves and only select those with a Creative Commons license. In contrast, most recent datasets (including Panda-70M chen2024panda , LVD-2M xiong2024lvd , Koala-36M wang2024koala , VidGen-1M tan2024vidgen , and OpenVid-1M nan2024openvid ) directly source data from HD-VILA-100M xue2022advancing . As a result: (1) they do not contribute new data but rather introduce a new data processing pipeline; (2) they must adhere to the same license as HD-VILA-100M, i.e., the Research Use of Data Agreement (R-UDA), which restricts commercial use. In addition, one of the most widely used datasets, WebVid-10M bain2021frozen , has been retracted due to potential copyright infringement.
  • VideoUFO inherits best practices from other curated datasets. We observe that recent datasets (1) feature detailed captions generated by multimodal large language models (MLLMs), (2) contain millions of video clips, and (3) are high-resolution (i.e., 720p). Therefore, to provide a large-scale and high-quality resource, our VideoUFO inherits these features.
Figure 5: The number of user-focused topics covered by recent video datasets. None successfully includes all user-focused topics.
4.2 Topics Coverage
This section analyzes the differences in topics coverage between the VideoUFO and existing datasets.
Calculation process. We calculate the number of user-focused topics covered by the existing datasets as follows:
  • Extract topics of recent video datasets. We repeat VidProM’s topic extraction process (Step 1 in Section 3) on the video datasets listed in Table 1. The number of topics extracted for each dataset is shown in Table 1 (##\##Topics).
  • Match the extracted topics with users’ focused ones. We observe that the same or similar topic may be described using different words. For example, both cathedral and church refer to a similar topic. Therefore, when comparing user-focused topics with those from each existing dataset, we choose semantic matching rather than a word-to-word approach. Specifically, (1) we first use $\mathtt{SentenceTransformers}$ reimers2019sentencebert to embed the two lists of topics; and (2) then for each user-focused topic, if there exists a topic in the existing dataset with a cosine similarity greater than $0.6$, we consider that user-focused topic to be covered by the existing dataset. Note that the threshold of $0.6$ is an empirical value, as we observe that most similar topics exhibit a cosine similarity greater than $0.6$.
Observations. The experimental results in Fig. 5 shows:
  • None of the existing datasets cover all user-focused topics in VideoUFO. Specifically, the most comprehensive dataset, InternVid wanginternvid , covers $821$ topics, accounting for $63.6\%$ of all user-focused topics. Furthermore, if we combine all existing datasets in Table 1 into the “largest” one, the coverage of user-focused topics will increase to $79.5\%$. However, this will lead to a higher computational burden when training text-to-video models.
  • Interestingly, a dataset may cover more user-focused topics than the actual number of topics it contains. For instance, OpenVid-1M nan2024openvid contains $671$ topics but covers $703$ user-focused ones. This is reasonable because two topics in our dataset, VideoUFO, may be similar and thus mapped to a single topic in OpenVid-1M nan2024openvid . Nevertheless, this does not compromise the quality of our VideoUFO since it only leads to some relative duplicated topics, while the videos in VideoUFO still accurately reflect real users’ focus and interests.
Limitation. The above topic analysis is based directly on the captions/descriptions provided by each recent video dataset. This is because downloading and re-captioning all these videos is prohibitively expensive in terms of network bandwidth and computation. The quality of these captions/descriptions may affect the number of extracted topics, as some may not be informative. To address this limitation, in the next section, we demonstrate that text-to-video models trained on these datasets fail to generate satisfactory videos for some topics.
5 VideoUFO Benefits Video Generation
Figure 6: The calculation process of BenchUFO. It is designed to evaluate whether a text-to-video model can effectively generate videos that contain user-focused topics. It comprises 791 concrete noun topics, each paired with $10$ real-world user-provided prompts.
5.1 Benchmark
This section details the proposed benchmark, BenchUFO, for evaluating text-to-video models’ performance on user-focused topics. We first introduce the design of the benchmark and then explain why it is reasonable.
Construction. As shown in Fig. 6, the calculation process of our BenchUFO includes: (1) Selecting prompts. First, we select 791 concrete nouns from $1,291$ user-focused topics. (Please refer to the next section for the rationale behind selecting concrete nouns.) Then, for each chosen topic, we randomly select $10$ text prompts from VidProM wang2024vidprom . All these prompts constitute the prompt set for our benchmark. (2) Generating videos. For each prompt, we use a text-to-video model to generate a corresponding video. (3) Describing videos. For each video, we use a multimodal large language model (here, we choose $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ Qwen2VL ) to understand and describe it. The instruction prompt is provided in the Appendix (Section B). (4) Calculating similarity. We use a sentence embedding model (in this case, $\mathtt{SentenceTransformers}$ reimers2019sentencebert ) to encode both the input prompt and the output description, and then compute the cosine similarity between them. We calculate the average similarity across $10$ prompts for each topic. A higher similarity score indicates better performance on that specific topic. Our benchmark computes and considers the $10$/$50$ worst-performing and $10$/$50$ best-performing topics.
Justification. This section provides justification for the proposed BenchUFO from three perspectives:
  • Why choose concrete nouns? This is due to the inconsistency between the input prompt and the output description when using abstract nouns. For instance, for the abstract noun topic “freedom” and the prompt “In a freedom world, ideal and love both exist”, a model may generate videos containing “dove”, “star”, or “heart”, and the video understanding model will summarize them as “a dove is playing…”, “a star is shining…”, and “a heart is beating…”. However, the embeddings of these descriptions are not expected to be close to those of the given prompt. This might create a false impression that the model cannot effectively generate content for that topic.
  • Why use video understanding model reasonable? This is because the video understanding model has advanced significantly with the advent of large language models, and it is reasonable to assume that these models have been exposed to images/videos containing user-focused topics. Here, we choose $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ Qwen2VL as the video understanding model. Trained on $800$ billion tokens of visual-related data and $600$ billion tokens of text data, it is expected to accurately comprehend these user-focused topics and faithfully summarize a video.
  • Why not use established benchmarks? This is because the existing benchmarks fail to reflect real-world scenarios.. Specifically, we note that there exists a well-established benchmark, VBench huang2024vbench , whose “object class” and “multiple objects” dimensions are similar to our BenchUFO. However, the prompts in VBench have two main drawbacks: (1) They cover only a limited number of topics/objects – specifically, just $79$. (2) They focus exclusively on common topics/objects (e.g., “bicycle”, “car”, and “airplane”), while many topics that users care about, such as “forest”, “sunset”, and “beach”, are missing.
Figure 7: The performance of both publicly available text-to-video models and our trained models on the proposed BenchUFO. The publicly available models are trained on various datasets, including both public and private ones. MVDiT nan2024openvid is trained on VidGen tan2024vidgen , OpenVid nan2024openvid , and VideoUFO, respectively. “Low/Top $N$” denotes the average score of the worst/best-performing $N$ topics.
  Models
Low $10$
Low $50$
Top $50$
Top $10$
Mira jumiradata $0.236$ $0.282$ $0.508$ $0.550$
Show-1 zhang2024show $0.266$ $0.303$ $0.524$ $0.564$
LTX-Video ltxvideo $0.268$ $0.310$ $0.532$ $0.574$
Open-Sora-Plan lin2024open $0.314$ $0.361$ $0.559$ $0.598$
TF-T2V wang2024recipe $0.316$ $0.359$ $0.560$ $0.595$
Mochi-1 genmo2024mochi $0.323$ $0.367$ $0.580$ $0.616$
HiGen qing2024hierarchical $0.352$ $0.394$ $0.589$ $0.625$
Open-Sora opensora $0.363$ $0.409$ $0.601$ $0.639$
Pika PikaLabs $0.365$ $0.404$ $0.583$ $0.619$
RepVideo si2025repvideo $0.368$ $0.402$ $0.589$ $0.619$
T2V-Zero khachatryan2023text2video $0.375$ $0.410$ $0.586$ $0.616$
CogVideoX yang2024cogvideox $0.383$ $0.419$ $0.601$ $0.629$
Latte-1 ma2024latte $0.384$ $0.421$ $0.592$ $0.636$
HunyuanVideo kong2024hunyuanvideo $0.388$ $0.427$ $0.612$ $0.645$
LaVie wang2024lavie $0.399$ $0.426$ $0.595$ $0.632$
Pyramidal jin2024pyramidal $0.400$ $0.433$ $0.606$ $0.647$
MVDiT-VidGen tan2024vidgen $0.382$ $0.426$ $0.594$ $0.626$
MVDiT-OpenVid nan2024openvid $0.401$ $0.437$ $0.609$ $0.645$
MVDiT-VideoUFO $\mathbf{0.442}$ $\mathbf{0.465}$ $\mathbf{0.619}$ $\mathbf{0.651}$
5.2 Observations
In this section, we evaluate current text-to-video models on this new benchmark and show that, with the help of VideoUFO, we achieve state-of-the-art performance. The quantitative and qualitative results are shown in Table 7 and Fig. 8, respectively. We observe that:
  • Current text-to-video models do not consistently perform well across all user-focused topics. Specifically, there is a score difference ranging from $0.233$ to $0.314$ between the top-$10$ and low-$10$ topics. These models may not effectively understand topics such as “giant squid”, “animal cell”, “Van Gogh”, and “ancient Egyptian” due to insufficient training on such videos.
  • Current text-to-video models show a certain degree of consistency in their best-performing topics. We discover that most text-to-video models excel at generating videos on animal-related topics, such as ‘seagull’, ‘panda’, ‘dolphin’, ‘camel’, and ‘owl’. We infer that this is partly due to a bias towards animals in current video datasets.
  • The proposed VideoUFO helps reduce the gap between the worst-performing and best-performing topics. To demonstrate the effectiveness of the proposed VideoUFO, we train an MVDiT nan2024openvid solely on VideoUFO. We find that the trained model achieves the highest low-$10$ scores (a $+4.2\%$ improvement compared to the current state-of-the-art) while maintaining performance on the top-$10$ topics. Visually, the trained model successfully generates videos on topics that other models previously could not.
  • The proposed VideoUFO outperforms other similar-scale datasets, such as OpenVid nan2024openvid and VidGen tan2024vidgen . To demonstrate that the improvement originates from our VideoUFO rather than from the MVDiT architecture nan2024openvid , we replace our VideoUFO with OpenVid nan2024openvid and VidGen tan2024vidgen . We find that when using the popular OpenVid nan2024openvid , the low-$10$ scores are similar to previous state-of-the-art models ($0.401$ vs. $0.400$). Furthermore, when training on VidGen tan2024vidgen , the performance is even lower.
Limitation. Due to limited computational resources, when validating the effectiveness of our dataset, we currently follow the video generation method from the academic community introduced by nan2024openvid , which already requires $32$ A100 GPUs. However, the quality of videos generated by this baseline remains limited. In the future, industry researchers could incorporate our dataset into their large-scale foundational model training pipelines to further explore its potential.
Figure 8: Visual comparisons between our approach (MVDiT-VideoUFO) and other text-to-video models. The model trained on VideoUFO outperforms the alternatives in generating user-focused topics. In the prompts, concepts in green indicate those successfully generated solely by our method, whereas those in yellow denote concepts successfully generated by both our model and the competing models. All generated videos are provided in the supplementary materials.
6 Ablation Studies on Curation of VideoUFO
This section performs ablation studies to investigate key components in curating VideoUFO, i.e., examining (1) the impact of the number of clusters in user-focused topic analysis and (2) the methods for clip verification. For more details, please refer to Appendix (Section C).
7 Conclusion
This paper presents a newly collected million-scale dataset for text-to-video generation with a focus on user needs. Beyond this, our dataset exhibits minimal overlap with existing datasets and is released under a more permissive license (i.e., CC BY). We first describe the curation process, which involves analyzing user-focused topics, searching for these topics, segmenting the resulting videos, and applying various post-processing techniques. Then, we highlight the differences between our dataset and existing ones in terms of fundamental attributes and topics coverage. Finally, we build a new benchmark to evaluate text-to-video models on user-focused topics and demonstrate that our dataset helps improve model performance in this regard. We encourage both the research community and industry to use our dataset to further advance the field of text-to-video generation.
Acknowledgments
We sincerely thank OpenAI for their support through the Researcher Access Program. Without their generous contribution, this work would not have been possible.
Appendix A The Prompt for Concluding Topics
Prompt: Could you describe the topic in the following short sentences using only 1-2 words? Please return only the topic (1-2 word) as a singular noun or in the base form of the verb.
Appendix B The Prompt for Describing Videos
[…]
Appendix C Ablation Studies on Curation of VideoUFO
This section performs ablation studies to investigate key components in curating VideoUFO, i.e., examining (1) the impact of the number of clusters used in user-focused topic analysis and (2) the methods used for clip verification.
The number of clusters. As shown in Fig. 9, we vary the number of pre-set clusters from $100$ to $3,000$ when analyzing the focused topics of text-to-video users. We observe that initially, as the number of pre-set clusters increases, the number of resulting topics also increases; however, after reaching approximately $2,000$ clusters, this increase stops and begins to fluctuate. Therefore, we finally choose $2,000$ as the number of pre-set clusters. This is reasonable because the number of user-focused topics is limited, and an excessive number of pre-set clusters would eventually merge together. Beyond this, we also experiment with DBSCAN and HDBSCAN [46] to automatically determine the number of clusters; however, we do not obtain reasonable results.
The method for clip verification. We randomly select $1,000$ video clips to conduct ablation studies on the methods used to verify whether a clip contains a specific topic. Feeding the entire video into GPT-4o (denoted as “Direct GPT-4o”) is considered the most accurate approach, and we compare the overlap of other cost-effective methods with it. As shown in Table 2, we observe that: (1) Feeding videos directly into GPT-4o is very expensive – processing $1,000$ clips costs \$$\$162.23$. (2) Using detailed captions instead of video clips reduces the API cost by approximately 5,400×5,400$\times$5 , 400 ×, while maintaining about $80\%$ prediction overlap. (3) Although using brief captions is even cheaper, the overlap drops by $14.9\%$, which is reasonable since a brief caption cannot capture all the information in a clip. (4) Interestingly, we find that GPT-4o-mini achieves slightly better performance than GPT-4o. In conclusion, for verifying clips, we choose detailed captions with GPT-4o-mini.
Figure 10: The statistical information of captions and video clips in the proposed VideoUFO dataset. The average word count for brief and detailed captions is $\mathbf{13.8}$ and $\mathbf{155.5}$, respectively, while the average clip duration is $\mathbf{12.6}$ seconds.
Figure 11: The statistical distributions of scores assessed by these six video quality metrics, which have been aligned with human perception. Researchers can use these scores to filter videos according to their specific training needs.
Appendix D The Distributions of Caption Lengths and Clip Durations
Fig. 10 presents the statistical distributions of caption lengths (for both brief and detailed captions) and clip durations.
Appendix E The Distributions of Video Quality Score
Fig. 11 shows the statistical distributions of scores evaluated in terms of subject consistency, background consistency, motion smoothness, dynamic degree, aesthetic quality, and imaging quality.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We follow this common practice for writing a paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See the Limitation at the end of Sections 4 and 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This is not a theory paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All the information needed to reproduce the main experimental results are provided in the main paper as well as the Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The dataset is publicly available at https://huggingface.co/datasets/WenhaoWang/VideoUFO, and the code is publicly available at https://github.com/WangWenhao0716/BenchUFO.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental settings and details are provided in the main paper as well as the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: All the experimental results are significant and stable, and error bars are not reported because it would be too computationally expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We include the experiments compute resources in each related section.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The data used in this study is publicly available on YouTube, and we do not anticipate any negative/positive societal impacts resulting from the release of our dataset.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We only scrape data with explicit CC BY 4.0 License to avoid any potential copyright infringement.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all the assets used in our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Please see the README file in the released dataset.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve that.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve that.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We only use LLM for writing, editing, or formatting purposes.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.